Find the linearization L(x,y) of the function at each point. f(x.y) = x2 + y2 + 1 a. (3,3) b. (1,3) a. L(x,y)=
Answer #1

Here, frac, y) = x+yíti fac (x, y) = a Co²+ y²+1) los e DC - 220 toto ( ək=0 for any une xe D= chotre peso luncu 20e doc cuns[ (x,y) = f(3,3) + (x-3) fx(3,3) + (9-3) fy (3,3) 19+ (2-3) 6+ (9-3)(6) 19 + 6X-18+ 64-18 :- LOX,9) = 6x + 64-17 So, lineariz